# Disentangled Representation Learning for Parametric Partial Differential Equations

### v1

@@ 4 Experiments
We assess the performance of DisentangO across a variety of physics modeling and discovery datasets. In the following, we denote a DisentangO model with a latent dimension of $n$ as DisentangO-$n$. Our evaluation focuses on the following key aspects. Firstly, we demonstrate the capability of DisentangO in forward PDE learning, and compare its performance with a number of relevant baselines. In particular, we select three NO-related baselines, i.e., FNO (Li et al., 2020c), NIO (Molinaro et al., 2023) and MetaNO (Zhang et al., 2023), as well as two non-NO baselines, i.e., VAE (Kingma and Welling, 2013) and $\beta$-VAE (Higgins et al., 2017). Secondly, we showcase the merits of DisentangO in inverse modeling for interpretable physics discovery. Thirdly, we perform parametric studies on the associated disentanglement hyperparameters. Lastly, we provide interpretations of the disentangled parameters and point out future directions. Further details are provided in Appendix C.
@@ 4.1 Supervised forward and inverse PDE learning
Table 1: Test errors and the number of trainable parameters. Bold number highlights the best method.
Models #param Test errors
data latent supervision
DisentangO 697,383 1.65% 4.63%
MetaNO 295,842 1.59% -
NIO 709,260 - 15.16%
FNO 698,284 2.45% 14.55%
VAE 697,672 49.97% 16.34%
$\beta$-VAE 697,672 51.16% 16.47%
@@
As the number of latent factors is fixed for supervised learning in this example, we defer our ablation study to the second experiment of semi-supervised learning. We report in Table 1 our experimental results from DisentangO as well as its comparison with the selected baselines. Note that only FNO, VAE and $\beta$-VAE are able to handle both forward and inverse problems, whereas MetaNO and NIO can only solve the forward problem and the inverse problem, respectively. With this caveat in mind, MetaNO achieves the best performance in forward PDE learning, with DisentangO performing comparably well and beating the third best model by 32.7% in accuracy. In terms of inverse modeling with full latent supervision, DisentangO is the only method that can hold the error below 5%, outperforming the second best method by 68.2%.
@@ 4.2 Semi-supervised mechanical MNIST benchmark
We next explore in the context of semi-supervised learning and apply DisentangO and relevant baselines to the benchmark Mechanical MNIST (MMNIST) dataset (Lejeune, 2020). MMNIST comprises 70,000 heterogeneous material specimens undergoing large deformation, each governed by a material model of the Neo-Hookean type with a varying modulus converted from the MNIST bitmap images. In our experiment, we take 500 images (with a 420/40/40 split for training/validation/test) and generate 200 loading/response data pairs per sample on a $29\times 29$ structured grid, simulating uniaxial extension, shear, equibiaxial extension, and confined compression load scenarios. Since only partial knowledge is available for each image (i.e., the corresponding digit), we apply semi-supervised learning to the latent factors to classify the digits.
Ablation study. Firstly, we investigate DisentangO’s predictability in forward PDE learning by comparing its performance to MetaNO (i.e., the base meta-learnt NO model without the hierarchical structure). As seen in Table 2, MetaNO achieves a forward prediction error of 2.68%, which serves as the optimal bound for DisentangO. As we increase the latent dimension in DisentangO from 2 to 15, the prediction error drops from 11.49% to 5.48% and converges to the optimal bound. Next, we study the role of the data loss term in disentanglement by gradually varying $\beta_{d}$ from $\beta_{d}=1$ to $\beta_{d}=1000$. In Table 2 we observe a consistent improvement in accuracy with the increase in $\beta_{d}$, where the boost in accuracy becomes marginal or slightly deteriorates beyond $\beta_{d}=100$. We thus choose $\beta_{d}=100$ as the best DisentangO model in this case. Besides offering an increased accuracy in forward modeling, the data loss term also enhances disentanglement as discussed in Section 3.2. This is evidenced by the illustration in Figure 2, where the unsupervised mutual information (MI) score that measures the amount of MI across latent factors consistently decreases in both DisentangO-2 and DisentangO-15 as we increase $\beta_{d}$ from $\beta_{d}=1$ (solid lines) to $\beta_{d}=100$ (dashed lines). On the contrary, the classification term poses a negative effect on disentanglement, as indicated by the increase in $\beta_{\text{cls}}$ leading to an increase in MI scores. This is reasonable because the classifier linearly combines all the latent factors to make a classification. The more accurate the classification, the stronger the correlation across latent factors. We then move on to study the effect of semi-supervised learning by comparing the model’s performance with and without latent partial supervision. While the models without semi-supervision is slightly more accurate in that the forward prediction accuracy with $\beta_{d}=1$ reaches 12.60% and 6.16% in DisentangO-2 and DisentangO-15, respectively, the latent scatterplot in Figure 3 reveals the inability of the DisentangO model without latent semi-supervision to acquire the partial knowledge of the embedded digits from data. In contrast, although the accuracy of DisentangO with latent semi-supervision slightly deteriorates due to the additional regularization effect from the classification loss term, it is able to correctly recognize the embedded digits and leverage this partial knowledge in disentangling meaningful latent factors.
Table 2: Test errors and the number of trainable parameters in MMNIST benchmark. Models are abbreviated (0=DisentangO-2, 1=DisentangO-5, 2=DisentangO-10, 3=DisentangO-15, 4=VAE, 5=$\beta$-VAE, 6=MetaNO). Bold number highlights the best method that can handle both forward and inverse settings.
Models 0 1 2 3 4 5 6
#param (M) 0.66 0.97 1.49 2.02 2.02 2.02 0.35
Error $\beta_{d}=1$ 12.82% 9.56% 7.36% 6.29% 16.34% 17.13% 2.68%
@@
Figure 3: MMNIST scatterplot with DisentangO-2 and $\beta_{d}=1$: left: $(\beta_{kl}=1,\beta_{cls}=100)$ (data test error 18.81%), middle: $(\beta_{kl}=10,\beta_{cls}=10)$ (data test error 16.94%), right: fully unsupervised DisentangO-2 without classification loss $(\beta_{kl}=100)$ (data test error 12.65%).
Comparison with additional baselines. We compare the performance of DisentangO with relevant additional baselines, i.e., VAE and $\beta$-VAE. We abandon the three NO-related baselines in this and the following experiments because they are black-box approximators and do not possess any mechanism to extract meaningful information without supervision. In this context, even the least accurate DisentangO-2 model (i.e., with $\beta_{d}=1$) with significantly fewer parameters outperforms the two selected baselines in Table 2 by 21.5% and 25.2%, respectively, with the best performing DisentangO-15 model beating the two baselines by 66.5% and 68.0% in accuracy, respectively. We do not proceed to compare the physics discovery capability between them as the baseline models are considerably inaccurate in forward prediction.
@@ 4.3 Unsupervised heterogeneous material learning
We now demonstrate the applicability of DisentangO in unsupervised learning in the context of learning synthetic human tissues that exhibit highly organized structures, with collagen fiber arrangements varying spatially. In this subject, it is critical to understand the underlying low-dimensional disentangled properties in the hidden latent space of this complex, high-dimensional microstructure, as inferred from experimental mechanical measurements. We generate two datasets, each containing 500 specimens and 100 loading/displacement pairs. The first dataset features variations in fiber orientation distributions using a Gaussian Random Field (GRF) and the second differs fiber angles in two segmented regions, separated by a centerline with a randomly rotated orientation.
Table 3: Test errors and the number of trainable parameters in experiment 3. Models are abbreviated (0=DisentangO-2, 1=DisentangO-5, 2=DisentangO-10, 3=DisentangO-15, 4=DisentangO-30, 5=VAE, 6=$\beta$-VAE, 7=MetaNO). Bold number highlights the best method that can handle both forward and inverse settings.
Models #param (M) Error
@@
0 0.63 26.33% 25.18%
1 0.94 17.51% 15.75%
2 1.46 10.76% 10.01%
3 1.98 7.11% 7.02%
4 3.55 5.33% 5.28%
5 3.55 61.10% -
6 3.55 57.04% -
7 0.32 2.67% -
@@ 5 Conclusion
We present DisentangO, a novel hyper-neural operator architecture designed to disentangle latent physical factors embedded within black-box neural operator parameters. DisentangO leverages a hierarchical neural operator architecture that extracts varying parameters of the governing PDE through a task-wise adaptive layer, then further disentangles these variations into distinct latent factors. By learning these disentangled representations, DisentangO not only enhances physical interpretability but also enables more robust generalization across diverse physical systems. We perform empirical evaluations across supervised, semi-supervised, and unsupervised learning contexts.
Limitations: Due to limited computational resource, our experiments are focused on learning a small-to-medium number ($<500$). Expanding the coverage to encompass a broader range of physical systems and enabling learning across more diverse parameters would be beneficial.
Acknowledgments and Disclosure of Funding
L. Zhang would like to acknowledge support by the AFOSR grant FA9550-22-1-0197. Y. Yu would like to acknowledge the support by the National Science Foundation under award DMS-2436624 and the support from the National Institute of Health under award 1R01GM157589-01. Portions of this research were conducted on Lehigh University’s Research Computing infrastructure partially supported by NSF Award 2019035.
Appendix A Proofs of identifiability analysis
@@
We proceed to derive the relation between $\bm{z}$ and $\hat{\bm{z}}$: since $\theta=r(\hat{\theta})$, with the invertibility assumption $\theta=\mathcal{H}^{-1}(\bm{z})$ and $\hat{\theta}=\hat{\mathcal{H}}^{-1}(\hat{\bm{z}})$, we obtain:
$\bm{z}=\mathcal{H}(\theta)=\mathcal{H}\circ r(\hat{\theta})=\mathcal{H}\circ r\circ\hat{\mathcal{H}}^{-1}(\hat{\bm{z}}).$
@@
$\sum_{i}\left(\dfrac{\partial^{2}q_{i}(z_{i},\bm{b})}{\partial z_{i}^{2}}\dfrac{\partial z_{i}}{\partial\hat{z}_{k}}\dfrac{\partial z_{i}}{\partial\hat{z}_{q}}+\dfrac{\partial q_{i}(z_{i},\bm{b})}{\partial z_{i}}\dfrac{\partial^{2}z_{i}}{\partial\hat{z}_{k}\partial\hat{z}_{q}}\right)+\dfrac{\partial^{2}\log|J_{h}|}{\partial\hat{z}_{k}\partial\hat{z}_{q}}=0\text{ .}$ (A.1)
@@
$\sum_{i=1}^{d_{z}}\left(\left(\dfrac{\partial^{2}q_{i}(z_{i},\bm{b}^{j})}{\partial z_{i}^{2}}-\dfrac{\partial^{2}q_{i}(z_{i},\bm{b}^{0})}{\partial z_{i}^{2}}\right)\dfrac{\partial z_{i}}{\partial\hat{z}_{k}}\dfrac{\partial z_{i}}{\partial\hat{z}_{q}}+\left(\dfrac{\partial q_{i}(z_{i},\bm{b}^{j})}{\partial z_{i}}-\dfrac{\partial q_{i}(z_{i},\bm{b}^{0})}{\partial z_{i}}\right)\dfrac{\partial^{2}z_{i}}{\partial\hat{z}_{k}\partial\hat{z}_{q}}\right)=0\text{ ,}$
@@ Appendix B Additional experimental details, results and discussion
Herein, we provide additional details in training and baseline models, as well as more results as a supplement of Section 4.
@@ B.1 Training details
In all experiments, we adopt the Adam optimizer for optimization and use a 16-layer DisentangO model. For fair comparison across different models, we tune the hyperparameters, including the learning rates, the decay rates, and the regularization parameters, to minimize the validation loss. Experiments are conducted on a single NVIDIA Tesla A100 GPU with 40 GB memory. A pseudo algorithm of all three scenarios is summarized in Algorithm 1.
@@
$L_{data}=$ $\dfrac{1}{S}\sum_{\eta=1}^{S}\left[-E(\log(p(\bm{u}|\bm{f},\theta^{\eta})))\right]\text{ ,}\quad L_{recon}=\dfrac{1}{S}\sum_{\eta=1}^{S}\left[-E_{q(\bm{z}^{\eta}|\theta^{\eta})}\log p(\theta^{\eta}|\bm{z}^{\eta})\right]\text{ ,}$
$L_{KL}=$ $\dfrac{1}{S}\sum_{\eta=1}^{S}\left[D_{KL}\left(q(\bm{z}^{\eta}|\theta^{\eta})||p(\bm{z}^{\eta})\right)\right]\text{ ,}\quad L_{semi}=L_{c}(\bm{z}^{\eta},c(\bm{b}^{\eta}))\text{ .}$
2:SC1: Unsupervised/SC3: Supervised

### v2

@@ 4 Experiments
Table 1: Test errors and the number of trainable parameters in experiment 1. Bold number highlights the best method.
Models #param per-epoch Test errors
time (s) data $\bm{z}$ (SC1)
DisentangO 697k 12.2 1.65% 4.63%
MetaNO 296k 9.8 1.59% -
NIO 709k 5.6 - 15.16%
FNO 698k 9.1 2.45% 14.55%
UFNO 720k 21.2 7.61% 11.23%
WNO 672k 183.5 8.09% 9.95%
PIANO 699k 16.9 7.99% 15.35%
CAMEL 654k 5.53 112.28% -
SMDP 671k 5.12 - 17.76%
InVAErt 707k 0.1 - 5.16%
FUSE 706k 2.2 - 4.99%
FUSE-f 707k 11.4 16.33% 6.19%
VAE 698k 2.8 49.97% 16.34%
convVAE 664k 2.8 81.11% 16.27%
$\beta$-VAE 698k 2.8 51.16% 16.47%
We assess DisentangO across various physics modeling and discovery datasets, denoting a model with latent dimension $n$ as DisentangO-$n$. Our evaluation focuses on several key aspects. Firstly, we demonstrate the capability of DisentangO in forward PDE learning, and compare its performance with a total of 14 relevant baselines. In particular, we select 8 NO baselines, i.e., FNO (Li et al., 2020c), UFNO (Wen et al., 2022), NIO (Molinaro et al., 2023), WNO (Tripura & Chakraborty, 2022), PIANO (Zhang et al., 2024), MetaNO (Zhang et al., 2023), FUSE (Lingsch et al., 2024) and its extension, as well as six non-NO baselines, i.e., CAMEL (Blanke & Lelarge, 2023), InVAErt (Tong et al., 2024), SMDP (Holzschuh et al., 2023), two VAE variants (Kingma & Welling, 2013) and $\beta$-VAE (Higgins et al., 2017). Secondly, we showcase the merits of DisentangO in inverse modeling for interpretable physics discovery. We perform parametric studies on the associated disentanglement parameters. Lastly, we provide interpretations of the disentangled parameters. Further details and an additional experiment for identifiability demonstration are provided in Appendices B and C.4, respectively.
@@ 4.1 Supervised Forward and Inverse PDE Learning
As the number of latent factors is fixed for supervised learning in this experiment, we defer our ablation study to the second experiment. We report in Table 1 our experimental results. Note that only FNO, PIANO, FUSE-f, VAE and its variations are able to handle both forward and inverse problems, whereas MetaNO can only solve the forward problem, and NIO, InVAErt, and FUSE can only solve the inverse problem. With this caveat in mind, MetaNO achieves the best performance in forward PDE learning, with DisentangO performing comparably well and beating the third best model by 32.7% in accuracy. In terms of inverse modeling with full latent supervision (SC1), DisentangO is the only method that can hold the error well below 5%, outperforming the second best method by 7.2% and the second best joint (i.e., simultaneous forward and inverse) solver by 25.2%.
@@ 4.2 Semi-Supervised Mechanical MNIST Benchmark
We consider semi-supervised learning and apply DisentangO to the Mechanical MNIST (MMNIST) benchmark (Lejeune, 2020). MMNIST comprises 70,000 heterogeneous material specimens undergoing large deformation, each governed by a material model of the Neo-Hookean type with a varying modulus converted from the MNIST bitmap images. In our experiment, we take 500 images (with a 420/40/40 split for training/validation/test) and generate 200 loading/response data pairs per sample on a $29\times 29$ structured grid, simulating uniaxial extension, shear, equibiaxial extension, and confined compression load scenarios. Since only partial knowledge is available for each image (i.e., the corresponding digit), we apply semi-supervised learning to the latent factors to classify the digits.
Table 2: Test errors and number of trainable parameters in experiment 2. DisentangO is abbreviated as DNO due to space limit. Bold number highlights the best method that can handle both forward and inverse settings.
Models DNO-2 DNO-5 DNO-10 DNO-15 VAE $\beta$-VAE MetaNO PIANO
#param (M) 0.66 0.97 1.49 2.02 2.02 2.02 0.35 2.05
$\beta_{d}=1$ 12.82% 9.56% 7.36% 6.29% 16.34% 17.13% 2.68% 13.73%
@@
Ablation study. Firstly, we investigate DisentangO’s predictability in forward PDE learning by comparing its performance to MetaNO (i.e., the base meta-learned NO model without the hypernetwork structure). As seen in Table 2, MetaNO achieves a forward prediction error of 2.68%, which serves as the optimal bound for DisentangO. As we increase the latent dimension in DisentangO from 2 to 15, the prediction error drops from 11.49% to 5.48% and converges to the optimal bound. Next, we study the role of the data loss term in disentanglement by gradually varying $\beta_{d}$ from $\beta_{d}=1$ to $\beta_{d}=1000$. In Table 2 we observe a consistent improvement in accuracy with the increase in $\beta_{d}$, where the boost in accuracy becomes marginal or slightly deteriorates beyond $\beta_{d}=100$. We thus choose $\beta_{d}=100$ as the best DisentangO model in this case. Besides offering an increased accuracy in forward modeling, the data loss term also enhances disentanglement as discussed in Section 3.2. This is evidenced by the illustration in Figure 3, where the unsupervised mutual information (MI) score that measures the amount of MI across latent factors consistently decreases in both DisentangO-2 and DisentangO-15 as we increase $\beta_{d}$ from $\beta_{d}=1$ (solid lines) to $\beta_{d}=100$ (dashed lines). On the contrary, the classification term poses a negative effect on disentanglement, as indicated by the increase in $\beta_{\text{cls}}$ leading to an increase in MI scores. This is reasonable because the classifier linearly combines all the latent factors to make a classification. The more accurate the classification, the stronger the correlation across latent factors. We then move on to study the effect of semi-supervised learning by comparing the model’s performance with and without latent partial supervision. While the models without semi-supervision is slightly more accurate in that the forward prediction accuracy with $\beta_{d}=1$ reaches 12.60% and 6.16% in DisentangO-2 and DisentangO-15, respectively, the latent scatterplot in Figure 2 reveals the inability of the DisentangO model without latent semi-supervision to acquire the partial knowledge of the embedded digits from data. In contrast, although the accuracy of DisentangO with latent semi-supervision slightly deteriorates due to the additional regularization effect from the classification loss term, it is able to correctly recognize the embedded digits and leverage this partial knowledge in disentangling meaningful latent factors.
Comparison with additional baselines. We compare DisentangO with additional baselines, i.e., VAE and $\beta$-VAE. We abandon the three NO-related baselines in this and the following experiments as they are black-box approximators and do not possess any mechanism to extract meaningful information without supervision. In this context, even the least accurate DisentangO-2 model (i.e., with $\beta_{d}=1$) with significantly fewer parameters outperforms the selected baselines in Table 2 by 21.5% and 25.2%, respectively, with the best performing DisentangO-15 model beating the baselines by 66.5% and 68.0% in accuracy, respectively. We do not proceed to compare the physics discovery capability between them as the baseline models are considerably inaccurate in forward prediction.
@@ 4.3 Unsupervised Heterogeneous Material Learning
Table 3: Test errors and number of trainable parameters in experiment 3. DisentangO is abbreviated as DNO due to space limit. Bold number highlights the best method for both forward and inverse settings.
Models #param (M) Test error
@@
DNO-2 0.63 26.33% 25.18%
DNO-5 0.94 17.51% 15.75%
DNO-10 1.46 10.76% 10.01%
DNO-15 1.98 7.11% 7.02%
DNO-30 3.55 5.33% 5.28%
VAE 3.55 61.10% -
$\beta$-VAE 3.55 57.04% -
MetaNO 0.32 2.67% -
PIANO 3.42 50.71%
We demonstrate DisentangO in unsupervised learning in the context of learning synthetic tissues that exhibit highly organized structures, with collagen fiber arrangements varying spatially. In this case, it is critical to understand the underlying low-dimensional disentangled properties in the hidden latent space of complex, high-dimensional microstructure, as inferred from experimental mechanical measurements. We generate two datasets, each containing 500 specimens and 100 loading/displacement pairs. The first dataset features variations in fiber orientation distributions using a Gaussian Random Field (GRF) and the second differs fiber angles in two segmented regions, separated by a centerline with a randomly rotated orientation.
@@ 5 Conclusion
We present DisentangO for disentangling latent physical factors embedded within black-box NO parameters. DisentangO leverages a hypernetwork-type NO architecture that extracts varying parameters of the governing PDE through a task-wise adaptive layer, and further disentangles these variations into distinct latent factors. By learning disentangled representations, DisentangO not only enhances physical interpretability but also enables robust generalization across diverse physical systems under different learning contexts. Limitations: Since the scalability of DisentangO is governed by the scalability of the NO backbone used for forward modeling, which has been extensively discussed in previous works, this study focuses on experiments involving high latent dimensions. Demonstrations on high-dimensional PDEs are beyond the scope of the current work.
Acknowledgments
L. Zhang was partially funded by the NSF grant DMS 2436624. Y. Yu would like to acknowledge the support by the National Institute of Health under award 1R01GM157589-01. Portions of this research were conducted on Lehigh University’s Research Computing infrastructure, partially supported by NSF Award 2019035.
Appendix A Identifiability analysis
A.1 Proof of the main theorems
@@
[…]
@@
[…]
@@
[…]
@@
A.2 Further discussion on the assumptions
Herein, we provide additional discussion on the validity and emprical validation for Assumptions 1-4.
As seen in the proof above, Assumptions 1 (Density Smoothness and Positivity) and 2 (Invertibility) are required to guarantee that there exists a smooth and injective mapping $h:=\mathcal{H}\circ r\circ\hat{\mathcal{H}}^{-1}$, from the ground-truth latent embedding $\bm{z}$ to the learned embedding $\hat{\bm{z}}$. Furthermore, the smoothness assumption further makes it feasible to take derivatives of $\bm{z}$ with respect to $\hat{\bm{z}}$, which supports the permutation-wise identifiability proof for Theorem 2. Here, we note that the smoothness assumption may possibly be relaxed to $C^{2}$. Assumptions 3 (Conditional Independence) and 4 (Linear Independence) are needed to show that the Jacobian of $h$ has one and only one non-zero component for each column. Without these assumptions, it is possible that the data from different $\mathbf{b}$ lack variability.
While the first three assumptions are common in many VAE architectures, the last assumption is plausible for many real-world data distributions. For instance, when the prior on the latent variables $p(\bm{z}|\bm{b})$ is conditionally factorial, where each element $z_{i}$ has a univariate exponential family distribution given conditioning variable $\bm{b}$:
$p(\bm{z}|\bm{b})=\prod_{i}\dfrac{Q_{i}(z_{i})}{Z_{i}(\bm{b})}\exp\left[\sum_{j=1}^{k}T_{i,j}(z_{i})\lambda_{i,j}(\bm{b})\right],$
where $Q_{i}$ is the base measure, $Z_{i}(\bm{b})$ is the normalizing constant, $T_{i,j}$ are the sufficient statistics, and $\lambda_{i,j}$ are the corresponding parameters depending on $\bm{b}$. This exponential family has universal approximation capabilities. Additionally, we note that this distribution is conditionally independent with
$q_{i}=\log(Q_{i}(z_{i}))-\log(Z_{i}(\bm{b}))+\left[\sum_{j=1}^{k}T_{i,j}(z_{i})\lambda_{i,j}(\bm{b})\right],$
and the linear independence indicates that the matrix formed by
$\omega(\bm{z},\bm{b}^{\eta})-\omega(\bm{z},\bm{b}^{0})=\left(\mathbf{T}^{\prime}(\lambda(\bm{b}^{\eta})-\lambda(\bm{b}^{0})),\mathbf{T}^{\prime\prime}(\lambda(\bm{b}^{\eta})-\lambda(\bm{b}^{0}))\right),\,\eta=1,\cdots,S$
has full rank $2d_{z}$.
In the fully supervised case, the conditional independence and linear independence assumptions are automatically guaranteed by picking proper distributions $p(\bm{z}|\bm{b})$ when designing the VAE architecture. In the semi-supervised and unsupervised cases, one can also validate these assumptions when the true values of $\bm{b}$ is given on some tasks, by inferring an empirical distribution of $p(\bm{z}|\bm{b})$ from these tasks. To investigate such a capability, in the additional synthetic experiment in Appendix C.4, we consider an unsupervised setting, estimate $p(\bm{z}|\bm{b})$ from the trained model, and check the linear independence condition by calculating the vector in equation 3.15 for each task $\bm{b}^{\eta}$ and forming an $S\times 2d_{z}$ matrix from all tasks. When the rank of this matrix is $2d_{z}$, it means that we can select $2d_{z}+1$ tasks from them with sufficient variability, such that the linear independence condition is satisfied. As a demonstration, in Appendix C.4 we validate this assumption on a synthetic dataset and show that the identifiability can be largely achieved with Gaussian distributions of distinct means and variances.
@@ Appendix B Additional experimental details, results and discussion
We provide additional details in training and baseline models, as well as more results as a supplement of Section 4.
@@ B.1 Training details
In all experiments, we adopt the Adam optimizer for optimization and use a $n$-layer DisentangO model, where $n=8$ in the first experiment and $n=16$ in the second and third experiments due to increased complexity. For fair comparison across different models, we tune the hyperparameters, including the learning rates, the decay rates, and the regularization parameters, to minimize the validation loss. Experiments are conducted on a single NVIDIA Tesla A100 GPU with 40 GB memory. A pseudo algorithm of all three scenarios is summarized in Algorithm 1.
@@
[…]
2:SC1: Supervised/SC3: Unsupervised